OpenStudy (anonymous):

if \[i ^{i ^{i ^{i ^{......^{\infty}}}}} \] = A+iB find A^2+ B^2

OpenStudy (anonymous):

@electrokid @shubhamsrg @phi @RadEn

OpenStudy (anonymous):

My Work Z = i^Z log Z = Z log(i) logZ = Z . i pie/2 .............and i am Stuck form here

OpenStudy (shubhamsrg):

log(x+iy) = ?

OpenStudy (anonymous):

log(x+iy) = log r + i theta

OpenStudy (anonymous):

r = |Z|

OpenStudy (shubhamsrg):

substitute z=x+iy and see if it helps r = sqrt(x^2 + y^2) theta = tan-1 (y/x) finally, compare real and imaginary parts

OpenStudy (anonymous):

Yup..Got it...THXXX